OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -7).

OpenStudy (anonymous):

Which way does the parabola opens?

OpenStudy (anonymous):

Down?

OpenStudy (nincompoop):

the focus should help reveal how the parabola opens

OpenStudy (anonymous):

good, so which general equation are you going to use?

OpenStudy (anonymous):

y=a(x-h)^+k

OpenStudy (anonymous):

you could, but it doesn't have a focus embedded in that equation. The better one is (x - h)^2 = 4p (y-k). Look familiar?

OpenStudy (anonymous):

Oh is the answer y = -1/28x2

OpenStudy (anonymous):

well, we'll see. Do you know what (h,k) and p stand for?

OpenStudy (anonymous):

(H,k) is the vertex, nd p idk

OpenStudy (anonymous):

p is the distance form the vertex to the focus. So what is that distance?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

technically it's 7. But because the parabola opens down, we know there will be a negative sign in front. So if you plug everything the general equation above, what do you get?

OpenStudy (anonymous):

I got -1/28x2

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

I just plugged in everything

OpenStudy (anonymous):

ok, so you did (x - 0)^2 = -4(7)(y-0) right?

OpenStudy (anonymous):

which then gave you y = -x^2/27 ?

OpenStudy (anonymous):

28*

OpenStudy (anonymous):

good